OpenStudy (anonymous):

put in exponetial form ^5sqrtsqrt3

OpenStudy (anonymous):

\[\sqrt[5]{\sqrt{3}}\]

OpenStudy (anonymous):

OK, well, consider this: \[\[\sqrt{x} = x^{1/2}\]\]

OpenStudy (anonymous):

you have squareroot of 3, then the 5th root of that value

OpenStudy (anonymous):

you can basically think of roots as the base raised to the fraction 1/the xth root...so \[\sqrt{x} =x^{1/2}\] \[\sqrt[y]{x}\] is \[x^{1/y}\]

OpenStudy (anonymous):

\[3^{5/2}\]?

OpenStudy (anonymous):

why 5 in the numerator?

OpenStudy (anonymous):

It was an educated guess

OpenStudy (anonymous):

the value of \[\sqrt[5]{\sqrt{3}}\] is (3^.5)^(1/5) is alot smaller than \[3^{5/2}\]

OpenStudy (anonymous):

so you have \[(3^{1/2})^{1/5}\]

OpenStudy (anonymous):

you can use the rules of exponents, and here you raising \[3^{1/2}\] to \[1/5\]

OpenStudy (anonymous):

so you would multiply the exponents here

OpenStudy (anonymous):

\[3^{1/10}\]?

OpenStudy (anonymous):

right on!

OpenStudy (anonymous):

haha thanks

OpenStudy (anonymous):

yeah, no problem, do you have anymore questions?

OpenStudy (anonymous):

wait, was this clear?

OpenStudy (anonymous):

Yeah, I'll have more exponetial questions lol. and yes it was explained clearly